# Motion Blender Gaussian Splatting for Dynamic Reconstruction

### v1

@@ Abstract
Gaussian splatting has emerged as a powerful tool for high-fidelity reconstruction of dynamic scenes. However, existing methods primarily rely on implicit motion representations, such as encoding motions into neural networks or per-Gaussian parameters, which makes it difficult to further manipulate the reconstructed motions. This lack of explicit controllability limits existing methods to replaying recorded motions only, which hinders a wider application. To address this, we propose Motion Blender Gaussian Splatting (MB-GS), a novel framework that uses motion graph as an explicit and sparse motion representation. The motion of graph links is propagated to individual Gaussians via dual quaternion skinning, with learnable weight painting functions determining the influence of each link. The motion graphs and 3D Gaussians are jointly optimized from input videos via differentiable rendering. Experiments show that MB-GS achieves state-of-the-art performance on the iPhone dataset while being competitive on HyperNeRF. Additionally, we demonstrate the application potential of our method in generating novel object motions and robot demonstrations through motion editing. Video demonstrations can be found at mlzxy.github.io/mbgs.
Figure 1: Our method reconstructs and renders dynamic scenes into 3D Gaussians and motion graphs from input videos. The learned motion graphs of a hand and a cat are visualized in the first and third rows. These graphs can be easily manipulated to create novel motions.
@@ 1 Introduction
Reconstructing and modeling dynamic 3D scenes is a fundamental challenge in computer vision, with applications spanning content creation, robotics, and augmented reality. Recent work on 3D Gaussian splatting has made significant progress in capturing dynamic scenes, enabling efficient and high-fidelity reconstruction [44, 17, 46, 24, 49, 47, 12].
@@
Existing Gaussian splatting-based methods utilize 3D Gaussians to represent geometry and appearance, paired with motion modules that determine the movements of Gaussians over time. For instance, 4DGaussians [46] and Deformable-GS [48] encode motion implicitly into neural networks. Shape-of-Motion [44] and STG [24] use shallower models that require dense per-Gaussian motion parameters. While these approaches achieve high-fidelity reconstruction of complex scenes, their implicit motion representation makes it difficult to manipulate the reconstructed motions. This lack of direct and explicit controllability restricts existing methods to replaying recorded motions only.
@@
Therefore, a key question to answer is: can we develop an explicit and sparse motion representation without compromising the ability to reconstruct complex scenes? To answer this question, we revisit some classical animation techniques [13]. Explicit motion representations, such as deformation graphs [42], harmonic coordinates [15], and cage [30], have been developed to animate objects with diverse motions. However, these classical methods focus on applying manually designed motions to mesh surfaces instead of 3D Gaussians [5]. Further, these methods are not able to reconstruct motion or geometry from visual data.
@@
Inspired by mesh animation techniques, we propose using kinematic trees and deformable graphs as motion representations for Gaussian splatting-based reconstruction. Kinematic trees are well-suited for capturing the motions of articulated objects, such as human bodies or robots, while deformable graphs, free from topological constraints, excel at modeling non-rigid object deformations. We collectively refer to kinematic trees and deformable graphs as motion graphs. Motions of graph links are propagated to individual Gaussians through dual quaternion skinning [16]. The influence of each graph link on a Gaussian is predicted by a learnable weight painting function. The graphs are initialized from point clouds and 2D keypoints at a canonical frame. Both the Gaussians and motion graphs are optimized end-to-end jointly from videos via differentiable rendering. We term our approach Motion Blender Gaussian Splatting (MB-GS). Our contributions can be summarized as follows.
  • We introduce a dynamic Gaussian splatting framework based on explicit and sparse motion graphs. Gaussian motions are predicted by blending the link motions.
  • We propose two types of parametric motion graphs—kinematic trees and deformable graphs—and a learnable weight painting function based on Gaussian kernels, along with optimization details in Sec. 3.4. Our method learns both 3D Gaussians and motion graphs jointly.
  • Compared with state-of-the-art, our method outperforms Shape-of-Motion on the iPhone dataset [8], and achieves competitive performance with 4DGaussians on HyperNerf [32]. Further, we demonstrate the application potential of our approach on creating novel object motions and synthesizing robot demonstrations through motion editing. This can lead to significant efficiency improvement in gathering training datasets for robot learning [43].
@@
Figure 2: Motion Blender Gaussian Splatting. Our framework explicitly represents motion using sparse graph structures. Static 3D Gaussians are associated with dynamic motion graphs by learnable weight painting. Then, link-wise motions are propagated to Gaussians through motion blending with dual quaternion skinning. We introduce two motion graph types: kinematic trees, ideal for capturing articulated structures like human bodies, and deformable graphs, designed for modeling non-rigid deformations in soft objects. The parameters of the motion graph, weight painting functions, and 3D Gaussians are jointly optimized end-to-end via differentiable rendering.
2 Related Works
@@ 3 Method
Sec. 3.1 introduces the overall motion blender framework for dynamic Gaussian splatting. Kinematic trees and deformable graphs, two types of motion graphs, are presented in Sec. 3.2. The weight painting function and the radiance Gaussian kernel are proposed in Sec. 3.3. Finally, the proposed optimization process is detailed in Sec. 3.4.
Preliminaries. Gaussian splatting represents a static 3D scene explicitly with a set $\mathcal{G}$ of 3D Gaussians. Each 3D Gaussian $g\in\mathcal{G}$ is parameterized by its pose $\mathbf{p}\in\operatorname{SE}(3)$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $o\in\mathbb{R}$ and color $\mathbf{c}\in\mathbb{R}^{3}$. 2D images can be rendered from $\mathcal{G}$ by blending the colors of overlapping Gaussians along the ray direction of each pixel.
Dynamic Gaussian splatting extends each Gaussian by making it time-dependent, $g_{t}$, where $t\in[0,T-1]$ and $T$ represents the number of frames in the source video. The common practice is to make the pose time-dependent, $\mathbf{p}_{t}$, and keep the rest static. $\mathbf{p}_{t}$ is often obtained from a model $f$ as in Eq. 1, where $\theta$ denotes the model’s parameters.
@@
Function $f$ is often implemented as a neural net such as a deformation field network [46, 48], or as a shallow model such as motion coefficients or polynomials [25, 24, 44] that require per-Gaussian motion parameters.
@@ 3.1 Motion Blender Gaussian Splatting
Our motion blender framework is designed based on two key ideas that set it apart from existing methods. 111 Represent motions with a sparse structure: Instead of associating dense motion parameters to each Gaussian (often numbering in the hundreds of thousands), we use a motion graph with far fewer parameters (often less than 100). The sparse link-wise motions are then propagated to each Gaussian through motion blending. 222 Represent motions explicitly: Instead of encoding motions implicitly into neural networks, our motion graph provides an explicit motion representation. This allows a straightforward visualization and manipulation of motions in the 3D space, such as transferring motion patterns from one object to a new object, or creating novel animations through motion graph editing.
Formally, we define the motion graph $G_{t}=(\mathcal{J}_{t},\mathcal{L})$ at time $t$ as a set of joints $\mathcal{J}_{t}$ with a set of edges $\mathcal{L}$, where $\max(|\mathcal{J}_{t}|,|\mathcal{L}|)\ll|\mathcal{G}|$. Each link $l\in\mathcal{L}$ is defined as $l=(s_{l},e_{l})$ and consists of the start and end joint indexes $s_{l},e_{l}\in[1,J]$. The positions of all the links at time $t$ are arranged as a matrix $X_{\mathcal{L},t}\in\mathbb{R}^{|\mathcal{L}|\times 6}$, where each row corresponds to the 3D coordinates of the start and end joints for a link in $\mathcal{L}$. Similarly, $\mathbf{x}_{t}\in\mathbb{R}^{3}$ denotes the position of a Gaussian with pose $\mathbf{p}_{t}$ at time $t$. We assume both the edges $\mathcal{L}$ and the number of joints $|\mathcal{J}_{t}|$ stay static, and only joint positions undergo motions over time. Let $P_{\mathcal{L},t}\in\operatorname{SE}(3)^{|\mathcal{L}|}$ be the poses of all the links at time $t$. Let $\mathbf{p}_{0}$ be the parameter that describes the initial pose of a Gaussian, then $\mathbf{p}_{t}$ is given by Eq. 2.
@@
$\mathcal{W}:\mathbb{R}^{3}\times\mathbb{R}^{L\times 6}\mapsto\Delta^{L-1}$ denotes the weight painting function, where $\Delta^{L-1}$ represents the probability vector space defined as $\{\mathbf{w}\in\mathbb{R}_{+}^{L}\ |\ \sum_{i=1}^{L}\mathbf{w}_{i}=1\}$. This function $\mathcal{W}$ estimates the influence of each link $l\in\mathcal{L}$ to a given Gaussian based on their relative positions at $t=0$.
$\mathcal{R}:\operatorname{SE}(3)^{L}\times\operatorname{SE}(3)^{L}\mapsto\operatorname{SE}(3)^{L}$ computes the relative rigid transformations between two sets of 3D poses. Specifically, $\mathcal{R}\left(P_{\mathcal{L},0},P_{\mathcal{L},t}\right)$ captures the link-wise SE(3) movements from time 00 to time $t$.
$\mathcal{B}:\operatorname{SE}(3)^{L}\times\Delta^{L-1}\mapsto\operatorname{SE}(3)$ represents the motion blending function, which computes the per-Gaussian motion by propagating the link-wise movements $\mathcal{R}\left(P_{\mathcal{L},0},P_{\mathcal{L},t}\right)$ to each Gaussian based on the weights assigned by $\mathcal{W}$ at $t=0$.
We implement $\mathcal{B}$ using dual quaternion skinning (DQS) [16]. DQS represents transformations using dual quaternions and computes weighted averages in a way that guarantees the resulting transformation remains valid in SE(3) space. The link-wise movements computed by $\mathcal{R}$ are analytically determined, as the relative SE(3) transformation can be straightforwardly derived. Therefore, our motion blender framework hinges on two key design choices: (1) Modeling graph motions by representing their temporal evolution through link poses $P_{\mathcal{L},t}$. (2) Defining a weight painting function $\mathcal{W}$ that accurately captures the influence of each graph link on a Gaussian’s motion. We address the former design choice by presenting two types of parametric motion graphs in Sec. 3.2. The latter, the design of $\mathcal{W}$, is detailed in Sec. 3.3. The overall motion blender gaussian splatting framework is visualized in Fig. 2.
@@ 3.2 Motion Graph Representation
To predict link poses $P_{\mathcal{L},t}$, we represent them as a parameterized function $\mathcal{P_{L}}(\theta,\phi)\in\operatorname{SE}(3)^{L}$. Here, $\theta$ encodes the graph parameters that determine the configuration of link poses at a given instant, and $\phi$ encodes the time-independent graph parameters. By modeling $\theta$ as a time-varying sequence $(\theta_{t})_{t=0}^{T-1}$, the link poses at time $t$ are defined as $P_{\mathcal{L},t}=\mathcal{P_{L}}(\theta_{t},\phi)$, where $\theta_{t}$ governs the kinematic state of the graph at time $t$. Two types of motion graphs, kinematic trees and deformable graph, are shown in Fig. 3.
@@
Figure 3: Motion Graphs. A kinematic tree (left) uses time-independent link lengths $\ell$ and dynamic joint rotations $\mathbf{r}_{t}\in\operatorname{SO}(3)$. Link poses (shown as colored coordinate axes) are computed through forward kinematics. A deformable graph (right) employs free-form topology parameterized by joint positions $\{\mathbf{n}_{i,t}\}$ and has non-rigid link deformations. Rigid per-link poses are obtained relative to each Gaussian position $\mathbf{x}_{0}$ and look-at transformation as in Eq. 3 and Eq. 4.
@@
Kinematic Tree. A kinematic tree is a hierarchical, acyclic graph with a root joint. This representation is ideal for capturing the kinematic chains of articulated objects such as human bodies, robot arms, or other multi-joint systems. We parameterize the tree using $\theta=(\mathbf{r},\mathbf{X}),\phi=(\ell_{i})_{i=1}^{L}$, where $\ell_{i}\in\mathbb{R}^{+}$ denotes the length of the $i$-th link, $\mathbf{r}\in\operatorname{SO}(3)^{J}$ represents the joint rotations, with $J$ denoting the number of joints, and $\mathbf{X}\in\operatorname{SE}(3)$ denotes the pose of the root joint. The $\mathcal{P_{L}}(\theta_{t},\phi)$ can be implemented via a forward kinematics algorithm [21], which propagates transformations from the root node through the tree, compounding rotations and translations across joints to derive every link pose. The forward kinematic procedure is differentiable, which enables the learning of $\theta$ and $\phi$ through back-propagation.
@@
Figure 4: Learned Motion Graphs and Weight Paintings. The first row overlays the learned motion graphs on the rendered images. The second row visualizes the painted weights (red) of the specific graph links (green).
@@
Deformable Graph. Unlike the kinematic tree, a deformable graph imposes no topological constraints, allowing joints to move freely in 3D space. This representation is ideal for modeling non-rigid deformations and surface variations in soft objects. We parameterize the graph as $\theta=(\mathbf{n}_{j}\in\mathbb{R}^{3})_{j=1}^{J},\phi=()$, where $\mathbf{n}_{j}$ denotes the 3D position of the $j$-th joint. However, the absence of topological constraints allows links between joints to stretch or compress, making it impossible to describe link poses $\mathcal{P_{L}}(\theta,\phi)$ using rigid transformations. To address this, we redefine the link pose as $P_{\mathcal{L}}(\theta,\mathbf{x})$, relative to a reference viewpoint $\mathbf{x}$. In practice, the position of each Gaussian is used as the reference viewpoint. For each link $l$ with start joint $\mathbf{n}_{s_{l},0}$ and end joint $\mathbf{n}_{e_{l},0}$ at $t=0$, we define $\mathbf{n}_{\mathbf{x}_{0},l}\in\mathbb{R}^{3}$ as the point on $l$ that has the minimal distance to $\mathbf{x}_{0}$. We refer $\mathbf{n}_{\mathbf{x}_{0},l}$ as the projection point at $t=0$. Next, we define the projection point $\mathbf{n}_{\mathbf{x}_{t},l}$ at time $t$ in Eq. 3.
@@
In other words, the projection point $\mathbf{n}_{\mathbf{x}_{t},l}$ moves proportionally to the stretching of link $l$. Next, the link pose $P_{\mathcal{L}}(\theta_{t},\mathbf{x}_{t})$ is derived via look-at transformation in Eq. 4.
@@
where $\operatorname{ray}(\mathbf{a},\mathbf{b})=\frac{\mathbf{b}-\mathbf{a}}{|\mathbf{b}-\mathbf{a}|}$ is the unit direction vector from $\mathbf{a}$ to $\mathbf{b}$. The look-at transformation $\mathscr{A}:\mathbb{R}^{3}\times S^{2}\mapsto\operatorname{SE}(3)$ maps the viewing position and direction to a $\operatorname{SE}(3)$ pose [40]. Note that look-at transformation requires defining an up-direction for each link. We organize the $L$ links into $\frac{L}{2}$ triangles and use their face normals as up-directions. Remark that there is a key distinction between our deformable graph and the deformation graph used in DynamicFusion [29]. DynamicFusion attaches SE(3) poses directly to joints only. Our method represents each joint as a $\mathbb{R}^{3}$ position, which simplifies motion-graph editing, and derives SE(3) link poses from joints. The link poses are then used to construct Gaussian motions.
Graph Connectivity $\mathcal{L}$ Initialization. The last problem to address is determining the graph connectivity $\mathcal{L}$, which defines the links between joints as integer pairs $l=(s_{l},e_{l})$, where $s_{l},e_{l}\in[1,J]$ denote the indices of the start and end joints. The connectivity $\mathcal{L}$ is treated as a fixed parameter, neither receiving gradients nor changing over time. For kinematic trees, connectivity is derived from domain-specific priors, such as lifting 2D human skeletons to 3D or extracting from robot models. For deformable graphs, we initialize $\mathcal{L}$ by randomly sampling and connecting points from a point cloud, using farthest point sampling [6] to ensure uniform coverage of the object’s surface, as illustrated in Fig. 6 and detailed in Sec. 3.4.

### v2

# Motion-Blender Gaussian Splatting for Dynamic Scene Reconstruction

@@ Abstract
Gaussian splatting has emerged as a powerful tool for high-fidelity reconstruction of dynamic scenes. However, existing methods primarily rely on implicit motion representations, such as encoding motions into neural networks or per-Gaussian parameters, which makes it difficult to further manipulate the reconstructed motions. This lack of explicit controllability limits existing methods to replaying recorded motions only, which hinders a wider application in robotics. To address this, we propose Motion Blender Gaussian Splatting (MBGS), a novel framework that uses motion graphs as an explicit and sparse motion representation. The motion of a graph’s links is propagated to individual Gaussians via dual quaternion skinning, with learnable weight painting functions that determine the influence of each link. The motion graphs and 3D Gaussians are jointly optimized from input videos via differentiable rendering. Experiments show that MBGS achieves state-of-the-art performance on the highly challenging iPhone dataset while being competitive on HyperNeRF. We demonstrate the application potential of our method in animating novel object poses, synthesizing real robot demonstrations, and predicting robot actions through visual planning. The source code, models, video demonstrations can be found at mlzxy.github.io/motion-blender-gs.
Keywords: Dynamic Scene Reconstruction, Gaussian Splatting
@@ 1 Introduction
Reconstructing and modeling dynamic 3D scenes is a fundamental challenge in computer vision, with applications on content creation, augmented reality, and robotics. Recent work on 3D Gaussian splatting has made significant progress in capturing dynamic scenes, enabling efficient and high-fidelity reconstruction [1, 2, 3, 4, 5, 6, 7]. Existing Gaussian splatting methods utilize 3D Gaussians to represent geometry and appearance, paired with motion modules that determine the movements of the Gaussians. For instance, 4D-Gaussians [3] and Deformable-GS [8] encode motion implicitly into neural networks. Shape-of-Motion [1] and STG [4] use shallower models that require dense per-Gaussian motion parameters. While these approaches achieve high-fidelity reconstruction, the reconstructed scenes cannot be easily altered or manipulated in simulation due to their implicit motion representation, which limits their use for robot manipulation planning. This lack of direct and explicit controllability restricts existing methods to simply replaying recorded motions.
@@
Therefore, a key question to answer is: can we develop an explicit and sparse motion representation without compromising the ability to reconstruct complex dynamic scenes? To answer this question, we revisit in this work some classical animation techniques [9]. Explicit motion representations, such as deformation graphs [10], harmonic coordinates [11], and cage [12], have been developed to animate objects with diverse motions. However, these classical hand-crafted methods focus on applying manually designed motions to mesh surfaces instead of differentiable models such as 3D Gaussians [13]. Further, these methods are not able to reconstruct motion or geometry from images.
Figure 1: Capabilities of Our Framework. Our method reconstructs and renders dynamic scenes into 3D Gaussians and motion graphs from input videos. The learned motion graphs for a hand and cat are shown with their corresponding rendered scenes (left). Our approach enables three key applications (right): ➊ Novel pose animation through motion graph editing, ➋ Robot demonstration synthesis by using robot kinematic chains as motion graphs, and ➌ Predicting robot actions by simulating graph movements to minimize the difference between rendered and goal images.
@@
Inspired by classical mesh animation techniques, we propose using kinematic trees and deformable graphs as motion representations for Gaussian splatting-based reconstruction. Kinematic trees are well-suited for capturing the motions of articulated objects, such as human bodies or robots, while deformable graphs, free from topological constraints, are ideal for modeling non-rigid object deformations. We collectively refer to both kinematic trees and deformable graphs as motion graphs. Motions of graph links are propagated to individual Gaussians through dual quaternion skinning [14]. The influence of each graph link on a Gaussian is predicted by a learnable weight painting function. The graphs are initialized from point clouds and 2D keypoints at a canonical frame. Both the Gaussians and motion graphs are optimized end-to-end jointly from videos via differentiable rendering. We term our approach Motion-Blender Gaussian Splatting (MBGS).
Our main contributions can be summarized as follows. (1) We introduce a new dynamic Gaussian splatting framework based on explicit and sparse motion graphs, which allows for straightforward robot manipulation planning in reconstructed scenes. Gaussian motions are predicted by blending the link motions. (2) We propose two types of parametric motion graphs—kinematic trees and deformable graphs—and a learnable weight painting function based on Gaussian kernels, along with optimization details in Sec. A.2. Our method learns both 3D Gaussians and motion graphs jointly. (3) Compared with state-of-the-art, our method outperforms Shape-of-Motion on the challenging iPhone dataset [15], and achieves competitive performance with 4DGaussians on HyperNerf [16]. Further, we demonstrate the applications of our approach on animating novel object motions, synthesizing real robot demonstrations, and predicting robot actions through visual planning. This can lead to significant efficiency improvement in gathering training datasets for robot learning [17].
@@
Figure 2: Motion Blender Gaussian Splatting. Our framework explicitly represents motion using sparse dynamic graphs. Static 3D Gaussians are associated with the graphs through learnable weight painting. Then, link-wise motions are propagated to the Gaussians through motion blending with dual quaternion skinning. We employ two motion graph types: kinematic trees, ideal for capturing articulated structures like human bodies, and deformable graphs, designed for modeling non-rigid deformations in soft objects. The parameters of the motion graph, weight painting functions, and 3D Gaussians are jointly optimized, end-to-end, via differentiable rendering.
@@ 2 Method
Preliminaries. Gaussian splatting represents a static 3D scene explicitly with a set $\mathcal{G}$ of 3D Gaussians. Each 3D Gaussian $g\in\mathcal{G}$ is parameterized by its pose $\mathbf{p}\in\operatorname{SE}(3)$, scale $\mathbf{s}\in\mathbb{R}^{3}$, opacity $o\in\mathbb{R}$ and color $\mathbf{c}\in\mathbb{R}^{3}$. 2D images can be rendered from $\mathcal{G}$ by blending the colors of overlapping Gaussians along the ray direction of each pixel. Dynamic Gaussian splatting extends each Gaussian by making it time-dependent, $g_{t}$, where $t\in[0,T-1]$ and $T$ represents the number of frames in the source video. The common practice is to make the pose time-dependent, $\mathbf{p}_{t}$, and keep the rest static. $\mathbf{p}_{t}$ is often obtained from a dynamic model $f$ as in Eq. 1, where $\theta$ denotes the model’s parameters.
@@
$f$ is often implemented as a neural net, e.g., a deformation field network [3, 8], or as a shallow model such as motion coefficients or polynomials [18, 4, 1] that require per-Gaussian motion parameters.
@@ 2.1 Motion Blender Gaussian Splatting
Our framework is designed based on two key ideas that set it apart from existing works. 111 Represent motions with a sparse structure: Instead of associating dense motion parameters to each Gaussian (often numbering in the hundreds of thousands), we use a motion graph with far fewer parameters (often less than 100). The sparse link-wise motions are then propagated to each Gaussian through motion blending. 222 Represent motions explicitly: Instead of encoding motions implicitly into neural networks, our motion graph is an explicit kinematic model. This allows a straightforward visualization and manipulation of motions in the 3D space, such as transferring motion patterns from one object to a new object, or creating novel animations through motion graph editing.
Formally, we define the motion graph $G_{t}=(\mathcal{J}_{t},\mathcal{L})$ at time $t$ as a set of joints $\mathcal{J}_{t}$ with a set of edges $\mathcal{L}$, where $\max(|\mathcal{J}_{t}|,|\mathcal{L}|)\ll|\mathcal{G}|$. Each link $l\in\mathcal{L}$ is defined as $l=(s_{l},e_{l})$ and consists of the start and end joint indexes $s_{l},e_{l}\in[1,|\mathcal{J}_{t}|]$. The positions of all the links at time $t$ are arranged as a matrix $X_{\mathcal{L},t}\in\mathbb{R}^{|\mathcal{L}|\times 6}$, where each row corresponds to the 3D coordinates of the start and end joints for a link in $\mathcal{L}$. Similarly, $\mathbf{x}_{t}\in\mathbb{R}^{3}$ denotes the position of a Gaussian with pose $\mathbf{p}_{t}$ at time $t$. We assume both the edges $\mathcal{L}$ and the number of joints $|\mathcal{J}_{t}|$ stay static, and only joint positions undergo motions over time. Let $P_{\mathcal{L},t}\in\operatorname{SE}(3)^{|\mathcal{L}|}$ be the poses of all the links at time $t$. Let $\mathbf{p}_{0}$ be the parameter that describes the initial pose of a Gaussian, then $\mathbf{p}_{t}$ is given by Eq. 2.
@@
$\mathcal{W}:\mathbb{R}^{3}\times\mathbb{R}^{|\mathcal{L}|\times 6}\mapsto\Delta^{|\mathcal{L}|-1}$ denotes the weight painting function, where $\Delta^{|\mathcal{L}|-1}$ represents the probability vector space defined as $\{\mathbf{w}\in\mathbb{R}_{+}^{|\mathcal{L}|}\ |\ \sum_{i=1}^{|\mathcal{L}|}\mathbf{w}_{i}=1\}$. This function $\mathcal{W}$ estimates the level of influence of each link $l\in\mathcal{L}$ on each Gaussian based on their relative positions at $t=0$.
$\mathcal{R}:\operatorname{SE}(3)^{|\mathcal{L}|}\times\operatorname{SE}(3)^{|\mathcal{L}|}\mapsto\operatorname{SE}(3)^{|\mathcal{L}|}$ returns relative rigid transforms between two sets of 3D poses. Specifically, $\mathcal{R}\left(P_{\mathcal{L},0},P_{\mathcal{L},t}\right)$ indicates the SE(3) movement from time 00 to time $t$ of each link in $\mathcal{L}$.
$\mathcal{B}:\operatorname{SE}(3)^{|\mathcal{L}|}\times\Delta^{|\mathcal{L}|-1}\mapsto\operatorname{SE}(3)$ represents the motion blending function, which computes the per-Gaussian motion by propagating the link-wise movements $\mathcal{R}\left(P_{\mathcal{L},0},P_{\mathcal{L},t}\right)$ to each Gaussian based on the weights assigned by $\mathcal{W}$ at $t=0$.
We implement $\mathcal{B}$ using dual quaternion skinning (DQS) [14]. DQS represents transformations using dual quaternions and computes weighted averages in a way that guarantees the resulting transformation remains valid in SE(3) space. The link-wise movements computed by $\mathcal{R}$ are analytically determined, as the relative SE(3) transforms can be straightforwardly derived. Therefore, our motion blender framework hinges on two key design choices: (1) Modeling graph motions by representing their temporal evolution through link poses $P_{\mathcal{L},t}$. (2) Defining a weight painting function $\mathcal{W}$ that accurately captures the influence of each graph link on a Gaussian’s motion. We address the former by presenting two types of parametric motion graphs in Sec. 2.2. The latter, the design of $\mathcal{W}$, is detailed in Sec. 2.3. The overall motion-blender Gaussian splatting framework is illustrated in Fig. 2.
@@ 2.2 Motion Graph Representation
To predict $P_{\mathcal{L},t}$, the link poses at time $t$, we use a parameterized function $\mathcal{P_{L}}(\theta,\phi)\in\operatorname{SE}(3)^{|\mathcal{L}|}$. Here, $\theta$ is the graph parameters that determine the link poses at a given instant, and $\phi$ is the time-independent graph parameters. By modeling $\theta$ as a time-varying sequence $(\theta_{t})_{t=0}^{T-1}$, the link poses at time $t$ are defined as $P_{\mathcal{L},t}=\mathcal{P_{L}}(\theta_{t},\phi)$, where $\theta_{t}$ describes the kinematic state of the graph at time $t$. Two types of motion graphs, kinematic trees and deformable graph, are shown in Fig. 3.
@@
Figure 3: Motion Graphs. A kinematic tree (left) uses time-independent link lengths $\ell$ and dynamic joint rotations $\mathbf{r}_{t}\in\operatorname{SO}(3)$. Link poses (shown as colored coordinate axes) in world coordinates are computed through forward kinematics. A deformable graph (right) employs free-form topology parameterized by joint positions $\{\mathbf{n}_{i,t}\}$ and has non-rigid link deformations. Rigid per-link poses are obtained relative to each Gaussian position $\mathbf{x}_{0}$ and look-at transformation as in Eq. 3 and Eq. 4.
@@
Kinematic Tree. A kinematic tree is a hierarchical, acyclic graph with a root joint. This representation is ideal for capturing the kinematic chains of articulated objects such as human bodies, robot arms, or other multi-joint systems. We parameterize the tree using $\theta=(\mathbf{r},\mathbf{X}),\phi=(\ell_{i})_{i=1}^{L}$, where $\ell_{i}\in\mathbb{R}^{+}$ denotes the length of the $i$-th link, $\mathbf{r}\in\operatorname{SO}(3)^{|\mathcal{J}_{t}|-1}$ represents the joint rotations, and $\mathbf{X}\in\operatorname{SE}(3)$ denotes the pose of the root joint. Then $\mathcal{P_{L}}(\theta_{t},\phi)$ can be implemented via a forward kinematics algorithm [19], which propagates transformations from the root node through the tree, compounding rotations and translations across joints to derive every link pose. The forward kinematic procedure is differentiable, which enables the learning of $\theta$ and $\phi$ through back-propagation.
@@
Figure 4: Learned Motion Graphs and Weight Paintings. The first row overlays the learned motion graphs on the images. The second row shows painted weights (red) of graph links (green).
@@
Deformable Graph. Unlike the kinematic tree, a deformable graph imposes no topological constraints, allowing the joints to move freely in the 3D space. This representation is ideal for modeling non-rigid deformations and surface variations in soft objects. We parameterize the graph as $\theta=(\mathbf{n}_{j}\in\mathbb{R}^{3})_{j=1}^{|\mathcal{J}_{t}|}$, where $\mathbf{n}_{j}$ denotes the 3D position of the $j$-th joint. However, the absence of topological constraints allows links between joints to stretch or compress, making it impossible to describe link poses using rigid transformations. To address this, we project each Gaussian $\mathbf{x}$ in the first frame on each link in the graph, and track the positions of the projected points in the remaining frames. Given each Gaussian $\mathbf{x}$, the matrix of the poses of all the links, denoted as $P_{\mathcal{L}}(\theta,\mathbf{x})$, is given by the 3D positions of the closest-point projections of $\mathbf{x}$ on the links in addition to the 3D directions of the links. Specifically, for each link $l$ with start joint $\mathbf{n}_{s_{l},0}$ and end joint $\mathbf{n}_{e_{l},0}$ at $t=0$, we define $\mathbf{n}_{\mathbf{x}_{0},l}\in\mathbb{R}^{3}$ as the point on $l$ that has the minimal distance to $\mathbf{x}_{0}$. We refer $\mathbf{n}_{\mathbf{x}_{0},l}$ as the projection point at $t=0$. Next, we define the projection point $\mathbf{n}_{\mathbf{x}_{t},l}$ at time $t$ in Eq. 3.
@@
In other words, the projection point $\mathbf{n}_{\mathbf{x}_{t},l}$ moves proportionally to the stretching of link $l$. Therefore, the link poses can be defined using the projection point $\mathbf{n}_{\mathbf{x}_{t},l}$, decoupling from non-rigid stretching deformations. Thus, the link pose $P_{\mathcal{L}}(\theta_{t},\mathbf{x}_{t})$ is derived via look-at transformation in Eq. 4.
@@
where $\operatorname{ray}(\mathbf{a},\mathbf{b})=\frac{\mathbf{b}-\mathbf{a}}{|\mathbf{b}-\mathbf{a}|}$ is the unit direction vector from $\mathbf{a}$ to $\mathbf{b}$. The look-at transformation $\mathscr{A}:\mathbb{R}^{3}\times S^{2}\mapsto\operatorname{SE}(3)$ maps the viewing position and direction to a $\operatorname{SE}(3)$ pose [20]. Note that look-at transformation requires defining an up-direction for each link. We organize the $|\mathcal{L}|$ links into $\frac{|\mathcal{L}|}{2}$ triangles and use their face normals as up-directions. Remark that there is a key distinction between our deformable graph and the deformation graph used in Dynamic Fusion [21]. Dynamic Fusion attaches SE(3) poses to joints only. Our method represents each joint as an $\mathbb{R}^{3}$ position and derives SE(3) poses at links, which simplifies motion-graph manipulation.
Graph Connectivity $\mathcal{L}$ Initialization. The last problem to address is determining the graph connectivity $\mathcal{L}$, which defines the links between joints as integer pairs $l=(s_{l},e_{l})$, where $s_{l},e_{l}\in[1,|\mathcal{J}_{t}|]$ denote the indices of the start and end joints. The connectivity $\mathcal{L}$ is treated as a fixed parameter, neither receiving gradients nor changing over time. For kinematic trees, connectivity is derived from domain-specific priors, such as lifting 2D human skeletons to 3D or extracting from robot models. For deformable graphs, we initialize $\mathcal{L}$ by randomly sampling and connecting points from a point cloud, using farthest point sampling [22] to ensure uniform coverage of the object’s surface, as illustrated in Fig. 6 and detailed in Appendix  A.2.